Internet Tools Summary

John December (decemj@rpi.edu)

March 11, 1994

11 Mar 94; Release 1.45

Copyright \begin{matrix}c \\ ◯ \end{matrix}1994 John December (decemj@rpi.edu). You can use this document for any personal or educational purpose. For-profit distribution requires my permission.

DISCLAIMER: Provided “as is” without expressed or implied warranty.

PURPOSE: to list tools available on the Internet that are used for network information retrieval (NIR), Computer-Mediated Communication (CMC), or other services. This is not meant to be a strict categorization or an exhaustive list, rather a reference catalog. I certainly would welcome comments and suggestions.

AUDIENCE: those getting started in understanding what you can do on the Internet in NIR and CMC; for those experienced, it collects and summarizes sources of information.

ASSUMPTION: you have access to and know how to use finger, ftp, gopher, http, telnet, email, or Usenet newsgroups.

REFERENCES: The references listed at the end of this document are very useful guides to these tools.

ACTION: The Action notation that I had previously described in this document is in the file internettools.tax.

NOTES:

1. Respect your access privileges to these tools.

2. This information changes; additions/comments welcome.

3. For demo purposes, I have used Unix commands; certainly Unix is not the only operating system required for these tools and forums. Apologies to those who don’t have Unix.

4. On the classification (NIR v. CMC): some tools could be used for either, such as telnet and email. I’ve placed them in what I feel are ”principle use” categories, e.g., telnet is used mostly for NIR, and email is commonly used for CMC (although both have applications in the other category).

5. Many of these tools have applicability off the Internet. Usenet, for example, is not confined to the Internet, and Internet email (and thus LISTSERV files) can be exchanged with communication systems off the Internet (BITNET, fidonet, commercial services.) So this list is not Internet (only) tools, but tools that can be used on the Internet.

Contents

1 HOW TO USE THIS DOCUMENT 1

2 NETWORK INFORMATION RETRIEVAL (NIR) TOOLS 3

3 COMPUTER-MEDIATED COMMUNICATION (CMC) FORUMS 7

4 PROTOCOLS/STANDARDS 10

5 REFERENCES 10

6 DESCRIPTION OF ITEMS 11

1 HOW TO USE THIS DOCUMENT 2

1 HOW TO USE THIS DOCUMENT

This document is formatted for LaTEX(for conversion to Postscript.) See the readme file for information on different formats for this document.

The tables in this document present information about the Internet. Tables of network-accessible

information are in the following form:

NETWORK ACCESSIBLE INFORMATION

ITEM NAME AM

PARAMETERS PARAMETERS

Where

• ITEM NAME = a short name describing the information or service.

• AM (ACCESS METHOD) = one of email, finger, ftp, gopher, http, phone, telnet, unix, Usenet news, wais.

• PARAMETERS = further access information depending on ACCESS METHOD:

ACCESS METHOD AND PARAMETERS

ACCESS METHOD PARAMETERS PARAMETERS
email  address  “letter-body”
finger  address  port #
ftp  host  path-to-directory/file-name
gopher  host
http host path-to-directory/file-name
news  newsgroup-name
phone  phone #
telnet host port # ‘login-name’
unix  command
wais  host  path

If any of the parameters are not on the line, they are not needed.

• All ftp access uses login-name ‘anonymous’ unless stated otherwise.

For information that is paper-based:

PAPER-BASED INFORMATION

Author year title

publisher or further information

INFORMATION ABOUT UPDATES OF THIS DOCUMENT

Internet-tools ftp ftp.rpi.edu pub/communications/internet-too ls.readme

Internet-tools Use ftp ftp.rpi.edu pub/communications/internet-too ls.use

Internet-tools Taxonomy ftp ftp.rpi.edu pub/communications/internet-too ls.tax

2 NETWORK INFORMATION RETRIEVAL (NIR) TOOLS 3

2 NETWORK INFORMATION RETRIEVAL (NIR) TOOLS

Utilities Finger

Definition: Finger retrieves information about user registered on a host computer. Finger description unix

man finger

Finger demo finger

decemj@rpi.edu

Finger Protocol ftp

nic.merit.edu documents/rfc/rfc1288.txt

Finger via email mail

b.liddicott@ic.ac.uk  Subject:  ”#finger

USER@HOST.DOM AIN”

Finger via gopher gopher rpi.edu

0decemj 79

Finger via telnet telnet rpi.edu

79

Finger via WWW http

cs.indiana.edu finger/gateway?USER

FingerInfo ftp

csd4.csd.uwm.edu pub/fingerinfo

Logfinger finger logfinger@twinbrook.cis.uab.edu Netfind

Definition: Netfind is a research prototype that provides a simple Internet ‘white pages’ user directory. Netfind description ftp

ftp.cs.colorado.edu pub/cs/distribs/netfind/README

Netfind paper ftp ftp.cs.colorado.edu pub/cs/techreports/schwartz/ASC

II/Netfind.Gathering.txt.Z

Netfind source ftp

ftp.cs.colorado.edu cs/distribs/netfind/

Netfind via http http alpha.acast.nova.edu

netfind.html

Netfind via telnet telnet mudhoney.micro.umn.edu

‘netfind’

Netfind via email mail listserv@brownvm.brown.edu “get netfind help”

Nslookup

Definition: Nslookup is an interactive program to query Internet domain name servers (gives IP address). Nslookup man page unix

man nslookup

Nslookup command unix nslookup domain-name

Ping

Definition: Ping requests echo from network host; see if remote host is up. Ping definition ftp

uxc.cso.uiuc.edu utils/ping/README

Ping via email mail

b.liddicott@ic.ac.uk Subject: ”#ping HOST@DOMAIN”

Ping demo unix /usr/etc/ping -s uwm.edu WHOIS

Definition: WHOIS provides info on registered network names. Whois description ftp

gopher.ucdavis.edu pub/IETF/WNILS/Discussion.Paper

Whois specs ftp

nic.merit.edu documents/rfc/rfc0954.txt

Whois Servers List ftp

sipb.mit.edu pub/whois/whois-servers.list

Whois via email mail

service@nic.ddn.mil Subject: whois HOST.DOMAIN

Whois via gopher gopher phantom.bsu.edu

4320

Whois via telnet telnet

rs.internic.net

Whois via Unix unix whois ”string”

X.500

Definition: X.500 (OSI Directory Service) provides globally distributed directory service. X.500 Definition ftp

nic.merit.edu documents/rfc/rfc1308.txt

X.500 Implement ftp

nic.merit.edu documents/fyi/fyi 11.txt

X.500 via telnet telnet

rs.internic.net

X.500 via gopher gopher barbarian.rs.itd.umich.edu 7777 Tools Archie

Definition: Archie locates files at anonymous FTP sites by filename search. Archie description ftp

archie.ans.net pub/archie/doc/whatis.archie

Archie manual page ftp

archie.ans.net pub/archie/doc/archie.man.txt

Archie via telnet telnet archie.ans.net

‘archie’

Archie via WWW http web.nexor.co.uk

archie.html

Archie via Unix unix

archie -s STRING

Archie via email mail archie@archie.unl.edu

“help”

Astra

Definition: Astra retrieve documents from databases. Astra Help mail astradb@icnucevm.bitnet

“help”

Astra documentation mail astradb@icnucevm.bitnet “GET META DOCUMENT”

Essence

Definition: Essence is a resource discovery system using indexes. Essence description ftp

ftp.cs.colorado.edu pub/cs/distribs/essence/README

Essence document ftp ftp.cs.colorado.edu pub/cs/distribs/essence/Essence

.txt.Z

File Service Protocol (Fsp)

Definition: Fsp is a conectionaless protocol for transferring files. FSP FAQ ftp

ftp.germany.eu.net pub/network/inet/fsp/FAQ

FSP INFO ftp

ftp.germany.eu.net pub/network/inet/fsp/INFO

FSP Sites mail

charro@ee.ualberta.ca Subject: fsp list “help”

FSP software ftp ftp.germany.eu.net pub/network/inet/fsp/ File Transfer Protocol (FTP)

Definition: FTP retrieves or puts copies of files at remote FTP sites. FTP how to ftp ftp.sura.net pub/nic/network.service.guides/

how.to.ftp.guide

FTP ftp

nic.merit.edu documents/rfc/rfc0959.txt

FTP FAQ ftp rtfm.mit.edu pub/usenet-by-group/news.answer

s/ftp-list/faq

FTP Interface http

hoohoo.ncsa.uiuc.edu ftp-interface.html

FTP via Unix unix

ftp

FTP via email mail ftpmail@decwrl.dec.com

“help”

FTP via telnet telnet

grind.isca.uiowa.edu

File Compression ftp

ftp.cso.uiuc.edu doc/pcnet/compression

FTP setup ftp cert.org

pub/tech tips/anonymous ftp

Jughead

Definition: Jughead gets menu information from gopher servers. Jughead description ftp boombox.micro.umn.edu pub/gopher/Unix/GopherTools/jug

head/jughead.ReadMe

Jughead Search gopher liberty.uc.wlu.edu

3002

Knowbot

Definition: Knowbot provides a uniform interface to heterogeneous remote directory services. Knowbot via telnet telnet info.cnri.reston.va.us

185

Maltshop

Definition: builds a veronica-access menu on your local gopher server. Maltshop script ftp veronica.scs.unr.edu veronica-code/

Netserv

Definition: Netserv is a server for access to data files and programs of general interest. Netserv mail netserv@frmop11.bitnet “get netserv helpfile”

Soft Pages

Definition: Soft Pages aids in retrieval of documents, software, and other resources from servers. Soft Pages documentation ftp ftp.tohoku.ac.jp

pub/spp

Project description ftp cs.ucl.ac.uk osi-ds/osi-ds-39-00.txt

Telnet

Definition: Telnet allows a user to login to a remote computer to use applications. Telnet how ftp ftp.sura.net pub/nic/network.service.guides/

how.to.telnet.guide

Telnet tips http

kufacts.cc.ukans.edu hytelnet html/TELNET.html

TELNET ftp

nic.merit.edu documents/rfc/rfc0959.txt

Telnet demo telnet downwind.sprl.umich.edu

3000

Telnet search gopher liberty.uc.wlu.edu

7 3004

Veronica

Definition: Veronica locates titles of Gopher items by keyword search. Veronica definition ftp

veronica.scs.unr.edu veronica/veronica-faq

Veronica FAQ ftp

veronica.scs.unr.edu veronica-docs/veronica-faq

Veronica Queries ftp

veronica.scs.unr.edu veronica-docs/how-to-query-vero nica

Gopher/Veronica documentation ftp

ftp.cso.uiuc.edu doc/net/uiucnet/vol6no1.txt

Veronica via gopher gopher veronica.scs.unr.edu Systems Alex

Definition: Alex provides transparent read of remote files at anonymous FTP sites. Alex description ftp

alex.sp.cs.cmu.edu usr0/anon/doc/NIR.Tool

Alex Web Page ftp

alex.sp.cs.cmu.edu usr0/anon/www/alex.html

Alex document ftp alex.sp.cs.cmu.edu usr0/anon/doc/intro.ps

Gopher

Definition: Gopher locates and retrieves resources using a graph of menus. Gopher guide ftp boombox.micro.umn.edu pub/gopher/docs/GopherGuide Jun 15b.txt

Gopher demo gopher

gopher.micro.umn.edu

Gopher via telnet telnet consultant.micro.umn.edu

‘gopher’

Gopher via email mail gophermail@ncc.go.jp

“help”

Gopher source ftp

boombox.micro.umn.edu pub/gopher/00README

Gopher documentation ftp

pit-manager.mit.edu pub/usenet/news.answers/gopher- faq

Gopher/Veronica documentation ftp

ftp.cso.uiuc.edu doc/net/uiucnet/vol6no1.txt

Gopher Newsgroup news comp.infosystems.gopher Moo Gopher

Definition: Moo Gopher is a MUD-based gopher interface. Moo Gopher blurb ftp

mailbase.ac.uk pub/lists/unite/files/jays hous e.txt

Moo Gopher via telnet telnet jh.ccs.neu.edu

1709

Prospero

Definition: Prospero provides user-centered view of remote files. Prospero document ftp prospero.isi.edu pub/prospero/doc/prospero-user-

manual.tex.Z

Prospero documentation ftp prospero.isi.edu pub/prospero/doc/README-prosper

o-documents

Wide Area Information Server (WAIS)

Definition: WAIS retrieves resources by searching indexes of databases. WAIS description ftp

quake.think.com wais/doc/overview.txt

WAIS ftp sunsite.unc.edu pub/docs/about-the-net/libsoft/

wais.txt

WAIS document ftp

quake.think.com wais/doc/wais-corp.txt

WAIS documentation ftp quake.think.com

wais/doc/

WAIS information at UWO ftp julian.uwo.ca

doc/wais/

WAIS, Free ftp ftp.cnidr.org pub/NIDR.tools/freeWAIS-0.1/REA DME

WAIS via telnet telnet

quake.think.com

WAIS via telnet telnet wais.com

‘wais’

WAIS via WWW http www.ncsa.uiuc.edu SDG/Software/Mosaic/Interfaces/

wais/wais-interface.html

WAIS newsgroup news

comp.infosystems.wais

WAIS clients ftp

sunsite.unc.edu pub/wais/clients/

freeWAIS ftp ls6-www.informatik.uni-dortmund .de

pub/wais/

WAISGATE http server.wais.com directory-of-servers.html

World Wide Web (WWW)

Definition: WWW retrieves resources by hyptertext browser of databases. WWW definition http info.cern.ch hypertext/WWW/LineMode/Defaults

/default.html

WWW ftp archive ftp

info.cern.ch pub/www/doc/

WWW guide http

www.hcc.hawaii.edu guide/www.guide.html

WWW Info http

www.bsdi.com server/doc/web-info.html

WWW General Info http golgi.harvard.edu

www.html

WWW 94 http

www1.cern.ch/WWW94/Welcome.html

WWW Research http

tdsrv1.ul.ie Research/WWW/www-research.html

WWW Book ftp emx.cc.utexas.edu pub/mnt/source/services/www/doc

/the www book.tex

WWW Growth http www.mit.edu

8001

WWW Bib http info.cern.ch hypertext/WWW/Bibliography.html 80

WWW faqs/guides http

cui www.unige.ch OSG/FAQ/www.html

WWW FAQ/gnat http

www.vuw.ac.nz non-local/gnat/www-faq.html

WWW FAQ/Boutell ftp

siva.cshl.org / boutell/www faq.html

WWW Kit http

www.geom.umn.edu docs/W3Kit/W3Kit.html

WWW Tools http info.cern.ch hypertext/WWW/Tools/Overview.ht ml

WWW MOMspider http

www.ics.uci.edu WebSoft/MOMspider/

WWW Clients http

info.cern.ch hypertext/WWW/Clients.html

WWW software http

info.cern.ch hypertext/WWW/Status.html

WWW Sites http

cui www.unige.ch OSG/Oscar/sites.html

WWW Getstats http www.eit.com

software/getstats/getstats.html

WWW Stats http

www.ics.uci.edu WebSoft/wwwstat/

WWW Usage Meter http siva.cshl.org

wusage.html

WWW via telnet/email http hypertext/WWW/FAQ/Bootstrap.htm l

WWW via telnet telnet

info.cern.ch

WWW via email mail listserv@info.cern.ch

“www URL”

WAISGATE http

server.wais.com directory-of-servers.html

WWW Newsgroup news comp.infosystems.www Interfaces Biomix

Definition: Biomix presents network resources as locations on screen maps. Note: Not known to be operational Biomix description ftp mailbase.ac.uk pub/lists/unite/files/biomix.tx t

Chimera

Definition: Chimera is a HTTP, FTP, Gopher client with an X/Athena interface. Chimera source ftp

ftp.cs.unlv.edu pub/chimera/

Chimera Home Page http www.unlv.edu chimera/index.html

Cello

Definition: Cello is a DOS-based Internet browser incorporating WWW, Gopher, FTP, Telnet, News. Cello source ftp ftp.law.cornell.edu pub/LII/Cello/

Compass

Definition: Compass is a system for Internet access to resources and tools for CERFnet subscribers; Compass description ftp

mailbase.ac.uk pub/lists/unite/files/compass.t xt Emacs World Wide Web Browser

Emacs WWW http

cs.indiana.edu elisp/w3/docs.html

Emacs WWW source ftp cs.indiana.edu pub/elisp/w3/

Fred

Definition: Fred is an interface to the OSI X.500 white pages directory service.

Fred telnet wp.psi.net ‘fred’

GN

Defintion: a multi-protocol (gopher0, http/1.0) interface GN Description ftp

ftp.acns.nwu.edu pub/gn/README

GN Source ftp ftp.acns.nwu.edu

pub/gn/

Hyper-G

Definition: Hyper-G is a large-scale, multi-user distributed hypermedia information system. Hyper-G via WWW http iicm.tu-graz.ac.at

ROOT 80

Hytelnet

Definition: Hytelnet provides Internet access info by hypertext (on network or TSR for DOS). Hytelnet description ftp

ftp.usask.ca pub/hytelnet/README

Hytelnet via WWW http

www.cc.ukans.edu hytelnet html/START.TXT.html

Hytelnet via telnet telnet info.ccit.arizona.edu

‘hytelnet’

Hytelnet search gopher liberty.uc.wlu.edu

3004

Mosaic for X/DOS/Mac

Definition: Mosiac is a client for network distributed hypermedia information and discovery. Mosaic home page http www.ncsa.uiuc.edu SDG/Software/Mosaic/NCSAMosaicH ome.html

Mosaic FAQ http www.ncsa.uiuc.edu SDG/Software/Mosaic/Docs/mosaic - faq.html

Mosaic ftp

ftp.ncsa.uiuc.edu Mosaic/README.Mosaic

Mosaic paper ftp

ftp.ncsa.uiuc.edu Mosaic/mosaic-papers/mosaic.ps. Z

Mosaic article http www.ugcs.caltech.edu

kluster/markoff.html

Amiga Mosaic http

insti.physics.sunysb.edu AMosaic/home.html

Mosaic unix mosaic

Lynx

Definition: an ASCII terminal browser for the World Wide Web. Lynx info ftp ftp2.cc.ukans.edu

pub/lynx/

Lynx Web info http www.cc.ukans.edu about lynx/www start.html

3 COMPUTER-MEDIATED COMMUNICATION (CMC) FORUMS 9

3 COMPUTER-MEDIATED COMMUNICATION (CMC) FORUMS

Interpersonal Email

Definition: Email allows a user to send message(s) to another user (or many users via mailing lists). Email How ftp ftp.sura.net pub/nic/network.service.guides/

how.to.email.guide

SMTP ftp

nic.merit.edu documents/rfc/rfc0821.txt

Email 101 ftp

mrcnext.cso.uiuc.edu etext/etext93/email025.txt

Finding Email addresses ftp sunsite.unc.edu pub/docs/about-the-net/libsoft/

email address.txt

College Email addresses ftp

pit-manager.mit.edu pub/usenet/soc.college/

InterNetwork Mail ftp

csd4.csd.uwm.edu pub/internetwork-mail-guide

Email info ftp ftp.uu.net

networking/mail/

Email echo mail

b.liddicott@ic.ac.uk Subject: ”#echo”

Email demo unix

mail user@host.domain

POPmail ftp boombox.micro.umn.edu pub/POPmail/

Talk

Definition: Talk provides real-time interactive text with another user. Talk man page unix

man talk

Talk demo unix talk user@host.domain Group Haven

Definition: A Haven is a network chat program that allows many people to talk to each other at once. Haven info ftp kidd.vet.purdue.edu

pub/haven/

Haven newsgroup news alt.internet.talk.haven

Listproc

Defintion: Listproc is a system that automates mailing lists and archives. Listproc info mail

listproc@stormking.com “help listproc”

Anastasios Kotsikonas mail tasos@cs.bu.edu

LISTSERV

Definition: LISTSERV is mailing-list server for group communication. LISTSERV definition mail

listserv@uacsc2.albany.edu “send listserv memo”

LISTSERV tips mail

listserv@bitnic.bitnet “get listserv tips”

LISTSERV searching mail

listserv@ulkyvm.bitnet “get database search”

LISTSERV managing mail

listserv@uhupvm1.uh.edu “get kovacs prv2n1”

LISTSERV tips mail

listserv@bitnic.bitnet “get listserv tips”

LISTSERV guide mail

listserv@earncc.bitnet “get lsvguide memo”

PAML http www.ii.uib.no

magnus/paml.html

Search LISTSERV http

alpha.acast.nova.edu cgi-bin/lists

Search LISTSERV WAIS wais munin.ub2.lu.se

210

Search LISTSERV gopher gopher dewey.lib.ncsu.edu 11/library/disciplines/library/ listgopher 70

Dartmouth Merged SIGL ftp dartcms1.dartmouth.edu

siglists/

Electronic Conferences ftp ksuvxa.kent.edu

library/

Search Academic Conferences wais munin.ub2.lu.se academic email conf 210

Majordomo

Definition: Majordomo is a mailing list manager. Majordomo mail Majordomo@GreatCircle.COM

“help”

Brent Chapman mail

brent@greatcircle.com

Majordomo Problems mail Majordomo-Owner@mv.mv.com

Moo

Definition: A Moo is an object-oriented Multiple User Dialogue (MUD). MediaMoo telnet purple-crayon.media.mit.edu

8888

Diversity University telnet erau.db.erau.edu

8888

Moo on a Web http jh.ccs.neu.edu

7043

Moo info gopher actlab.rtf.utexas.edu Multiple-User Dialogue (MUD)

Definition: MUDs are real-time interaction usually for social role-playing. Note: Variants include MUCK, MUSH, MUSE, MOO. MUD FAQ ftp

ftp.math.okstate.edu pub/muds/misc/mud-faq/

MUD Lists ftp caisr2.caisr.cwru.edu

pub/mud/

MUD WWW Page http math.okstate.edu

mud.html 8001

MUDS http

ftp.ccs.neu.edu pub/mud/www/muds.html

MUD types http

csugrad.cs.vt.edu soc/mud types.html

MUD info gopher

actlab.rtf.utexas.edu

MUD research gopher

actlab.rtf.utexas.edu

MUDWHO via WWW http

math.okstate.edu mudwho.html 8001

MUD/Lysators http

www.lysator.liu.se nobw/mud/main.html 7500

MUD Announcements news

rec.games.mud.announce

MUD Administration news

rec.games.mud.admin

MUD Miscellaneous news

rec.games.mud.misc

MUD Tiny news

rec.games.mud.tiny

MUD Diku news rec.games.mud.diku

Procmail

Definition: Procmail is a mail manager–create mail-servers, mailing lists, sort your incoming mail. Procmail source ftp

ftp.informatik.rwth-aachen.de pub/packages/procmail/

Procmail mailinglist mail procmailrequest@informatik.rwt h-aachen.de

Mass Interactive Webbing (IW)

Definition: Interactive Webbing gives people a common space for network distributed multimedia writing. Free For All http

south.ncsa.uiuc.edu Free.html 8241

BBS Home Page http indian.irsa.jrc.it BBS/home.html Internet Relay Chat (IRC)

Definition: IRC provides real-time, many-many text discussion divided into channels. IRC definition ftp

nic.merit.edu documents/rfc/rfc1459.txt

IRC Info/Larsson http

eru.dd.chalmers.se home/f88jl/Irc/ircdocs.html

IRC FAQ ftp cs.bu.edu

irc/support/alt-irc-faq

IRC guide ftp dorm.rutgers.edu pub/Internet.documents/irc.basi

c.guide

IRC documentation ftp cs.bu.edu

irc/support/IRCprimer1.1.txt

IRC Web Page http www.kei.com

irc.html

IRC via telnet telnet sci.dixie.edu

6677

IRC connections http

alpha.acast.nova.edu irc/connect.html

IRC via Unix unix

irc

IRC newsgroup news

alt.irc

IRC ircii news

alt.irc.ircii

IRC lamers news

alt.irc.lamers

IRC recovery news

alt.irc.recovery

IRC undernet news alt.irc.undernet Internet Talk Radio (ITR)

Definition: ITR is an audio multicast on the Internet. ITR introduction ftp

sunsite.unc.edu pub/talk-radio/ITRintro.readme

ITR FAQ mail

info@radio.com

ITR Web http

www.ncsa.uiuc.edu radio/radio.html

Internet Talk Radio sites mail

sites@radio.com

ITR FAQ ftp rftm.mit.edu pub/usenet/news.answers/interne

t-talk-radio/

ITR Newsgroup news alt.internet.talk-radio

Mbone

Definition: Mbone is a live audio and video multicast virtual network on top of Internet. Mbone FAQ (text) ftp

venera.isi.edu mbone/faq.txt

Mbone FAQ Web http

www.research.att.com mbone-faq.html

Mbone Tools http www.eit.com

techinfo/mbone/mbone.html

Usenet

Definition: USENET provides asynchronous text discussion on many topics separated into newsgroups. USENET description ftp rtfm.mit.edu pub/usenet/news.answers/what-is

-usenet/

USENET FAQs ftp rtfm.mit.edu pub/usenet/news.announce.newuse rs/

USENET documentation ftp

wuarchive.wustl.edu doc/misc/acn/acn4-5.txt.Z

NNTP ftp

nic.merit.edu documents/rfc/rfc0977.txt

Usenet Lists ftp

rtfm.mit.edu pub/usenet/news.groups/

Search Usenet WAIS wais munin.ub2.lu.se

210

Usenet filter http woodstock.stanford.edu

2000

Usenet filter mail netnews@db.stanford.edu

“help”

Beginners group news

news.announce.newusers

Usenet addicts news

alt.usenet.addict

Usenet humor discussion news

alt.humor.best-of-usenet.d

Usenet humor news

alt.humor.best-of-usenet

Usenet culture news

alt.culture.usenet

Usenet recovery news

alt.usenet.recovery

Usenet reader unix

rn newsgroup-name

Usenet guide http

ocf.berkeley.edu pub/trnint-3.3.html

Usenet post via email mail hierarchy-group-name@cs.utexas. edu Subject: Your Subject “Your Co ntents”

Usenet via WWW http info.cern.ch hypertext/DataSources/News/Grou

ps/Overview.html

Usenet via gopher finger

lesikar@tigger.stcloud.msus.edu

Usenet via telnet telnet sol.ctr.columbia.edu

119

Usenet periodics finger

nichol@stavanger.sgp.slb.com

Usenet posters mail mail-server@rtfm.MIT.EDU “send usenet-addresses/FIRST L AST”

4 PROTOCOLS/STANDARDS 12

4 PROTOCOLS/STANDARDS

INTERNET PROTOCOLS/SERVERS

IP ftp

nic.merit.edu documents/rfc/rfc0791.txt

Internet Standards ftp

nic.merit.edu documents/std/INDEX.std

Kerberos ftp

rtfm.mit.edu pub/usenet/comp.protocols.kerbe ros/

NFS ftp

rtfm.mit.edu pub/usenet/comp.protocols.nfs/

PPP ftp

rtfm.mit.edu pub/usenet/comp.protocols.ppp/

PPP FAQ http

cs.uni-bonn.de ppp/faq.html

SLIP ftp

vtucs.cc.vt.edu filebox/nyman/whatslip.txt

SNMP ftp

rtfm.mit.edu pub/usenet/comp.protocols.snmp/

SOCKS ftp ftp.nec.com pub/security/socks.cstc/What Is SOCKS.CSTC

TCP ftp

nic.merit.edu documents/rfc/rfc0793.txt

TCP-IP FAQ ftp

rtfm.mit.edu pub/usenet/comp.protocols.tcp-i p/

Winsock ftp

nebula.lib.vt.edu pub/windows/winsock/wtcpip03.as c

OTHER STANDARDS/PROTOCOLS

ACC ftp

research.att.com dist/anoncc/accinet.ps.Z

Telecom stds ftp rftm.mit.edu pub/usenet/news.answers/standar dsfaq

ISO Protocols news

comp.protocols.iso

ISO FAQ ftp

rtfm.mit.edu pub/usenet/comp.protocols.iso/

OSI ftp rftm.mit.edu pub/usenet/news.answers/osi-pro

tocols

URL guide http

www.ncsa.uiuc.edu demoweb/url-primer.html

URL syntax http info.cern.ch hypertext/WWW/Addressing/URL/5

BNF.html

Z39.50 W3 Page http

www.vtls.com market/z39/z39page.html

5 REFERENCES

OTHER STANDARDS/PROTOCOLS

Internet Tools EARN ftp ns.ripe.net

earn/earn-resource-tool-guide.t xt

Internet Tools NIR ftp

mailbase.ac.uk pub/lists/nir/files/nir.status. report

Internet Systems UNITE ftp

mailbase.ac.uk pub/lists/unite/files/systems-l ist.txt

Net Mgt Tools ftp

nic.merit.edu documents/fyi/fyi 02.txt

PC Communications http www.rtd.com pcnfsfaq/faq.html

6 DESCRIPTION OF ITEMS 13

6 DESCRIPTION OF ITEMS

Item Descriptions

Internet-tools describes Internet tools project and other formats of the information Internet-tools Use some tips and instructions on using this file Internet-tools Taxonomy

some notes toward a taxonomy of Internet tools

Finger description Unix manual page for finger Finger demo

find out about user decemj@rpi.edu

Finger Protocol The Finger User Information Protocol, by D. Zimmerman Finger via gopher gopher to HOST, port 79, path 0USER-ID Finger via telnet access finger via telnet HOST=host on which USER is located, then type USER Finger via WWW access finger information on USER via World Wide Web FingerInfo

a script to get information via finger, by Scott Yanoff

Logfinger

A Program to Log a Fingerer

Netfind description a way to find someone on the Internet Netfind via telnet using telnet to find someone on the Internet Netfind via email use email to find someone on the Internet Nslookup man page Unix man page Nslookup command command to find out about comand-name Ping definition description of Ping and its use Ping demo

check to see if host uwm.edu is up and roundtrip time

Whois description describes whois Whois specs

NICNAME/WHOIS

Whois Servers List List of Internet whois servers, by Matt Power Whois via telnet at the prompt, enter whois host-name Whois via Unix the Unix command for whois, where ‘string’ is the organization/person name or In-

ternet domain name

X.500 Definition rfc specifying X.500 X.500 via telnet at the prompt, enter X500whois host-name Archie description archie - An Electronic Directory Service for the Internet, by Deutsch, Emtage, and

Heelan

Archie manual page reference manual page for archie Archie via telnet a telnet demo of archie Archie via WWW Hypertext Archie Servers Archie via Unix Unix command line for hosts with installed archie client, search for STRING Archie via email use archie via email Astra Help

the help file for Astra

Astra documentation Astra User Guide Essence description The distribution includes the following files and directories Essence document Essence – A Resource Discovery System Based on Semantic File Indexing FTP how to

A guide to using ftp

FTP

File Transfer Protocol

FTP FAQ

Anonymous FTP Frequently Asked Questions (FAQ) List

FTP Interface look at ftp site lists FTP via Unix Unix command for ftp, once in, type help FTP via email get files at anonymous ftp sites via email. FTP via telnet site at Univ of IA File Compression File compression, archiving, and text¡-¿binary formats, by David Lemson Jughead description description of current status, by Rhett ‘Jonzy’ Jones Knowbot via telnet use a knowbot service via telnet; at the prompt, type ‘help’ Netserv

a server for access to data files and programs of general interest

Project description Soft Pages Project, by Johannsen and Mansfield Telnet how

a guide to using telnet

TELNET

Telnet Protocol

Telnet demo check the weather Telnet search find telnet connections based on keyword search Veronica definition frequently asked questions and answers about veronica Veronica via gopher Search ALL of Gopherspace using veronica Alex description NIR (Networked Information Retrieval) description of Alex Alex Web Page Alex FTP Filesystem Gopher guide Internet Gopher User’s Guide, ed. Paul Linder Gopher demo a session demonstrating gopher at Univ of Minnesota Gopher via email use a gopher via email Gopher source description of gopher client sources Moo Gopher blurb a short description of Moo gopher Prospero document Prospero user’s manual, by Neuman and Augart Prospero documentation

description of prospero documentation available

WAIS description an overview of WAIS WAIS

A Sketch of An Overview, by Jeff Kallem

WAIS, Free

a UNIX based, freely available information server.

WAIS clients

client to WAIS in different interfaces (X)

WAISGATE

WAIS ¡-¿ WWW gateway

WWW definition home page for WWW WWW ftp archive retrieves resources by hyptertext browser of databases WWW guide

Entering the WWW–A Guide to Cyberspace by Kevin Hughes

WWW 94 First International Conference on the World-Wide Web May 25-26-27 1994, CERN, Geneva (Switzerland)

WWW Research links to documents on W3 research WWW Book

a book describing the World Wide Web project

WWW Bib

Bibliography for the World Wide Web

WWW faqs/guides a list by Oscar Nierstrasz WWW FAQ/gnat gnat’s World Wide Web Frequently Asked Questions and Answers Lists WWW FAQ/Boutell FAQ on WWW (World Wide Web Frequently Asked Questions), by Thomas Boutell WWW Kit

a system for building interactive graphical applications for the World Wide Web

WWW Tools

Tools for WWW providers

WWW MOMspider A World-Wide Web Robot for Multi-Owner Maintenance WWW Clients a list of programs to allow you to access the WWW from your own computer WWW software list of products related to WWW WWW Getstats a versatile World-Wide Web server log analyzer. It takes the log file from your CERN,

NCSA, Plexus, GN, or UNIX Gopher server and spits back all sorts of statistics

WWW Stats

collect usage statistics and output a log summary in HTML forma

WWW Usage Meter collects information about usage of a WWW site, by Thomas Boutell WWW via telnet/email

list of sites

WWW via telnet using WWW via telnet WWW via email obtain a web file (e.g., http) via email; URL = Uniform resource locator WAISGATE

WAIS ¡-¿ WWW gateway

Biomix description BIOMIX current status, by Marcus Pattloch Chimera Home Page a HTTP, FTP, Gopher client with X/Athena interface, by John Kilburg

Cello source source code

Compass description Internet Compass-your network guide, CERFnet offers new service for easier

Internetting

Emacs WWW an Emacs subsystem that allows the user to browse the World Wide Web Fred

Performance Systems International White Pages Pilot Project

Hyper-G via WWW Hyper-G gateway at Graz Univ. of Technology Hytelnet description Peter Scott’s Hypertext database of publicly accessible Internet sites Hytelnet search given a keyword, returns hytelnet entries Mosaic home page describes Mosaic, Internet information browser and World Wide Web client Mosaic description of how to download and run NCSA Mosaic Mosaic paper NCSA Mosaic Technical Summary, by Marc Andreessen Mosaic article an HTML version of a New York Times article about Mosaic by John Markoff Mosaic

bring up Mosaic on system with client installed

Lynx Web info Welcome to Lynx and the world of the web Email How

Introduction to Sending and Receiving Electronic Mail

SMTP

Simple Mail Transfer Protocol

Email 101

describes how to use email as well as other Internet features, by John Goodwin

Finding Email addresses

how do I find someone’s email address?

College Email addresses

information on email addresses at colleges

InterNetwork Mail methods of sending mail from one network to another, by John Chew and Scott Yanoff Email info

repository of types of email interfaces and tools

Email echo

sends email back to you

Email demo

Unix (UCB) mail command

Talk man page Unix manual page for talk Talk demo

where user@host.domain is known

Haven info

a repository of documents, source code, logs and other information

Anastasios Kotsikonas Listproc creator LISTSERV definition the help file for listserv LISTSERV searching tips on searching a listserv discussion for information LISTSERV managing How to Start and Manage a BITNET LISTSERV Discussion Group, by Kovacs, Mc-

Carty, and Kovacs

LISTSERV tips List Management Tips for LISTSERV Postmasters and List Owners, by Lisa M. Covi LISTSERV guide The Listserv Guide for General Users PAML

Publicly Available Mailing Lists, by Stephanie da Silva

Search LISTSERV search for interesting email lists to join Search LISTSERV WAIS

WAIS-based search for a Usenet and LISTSERV conferences

Search LISTSERV gopher

searching and retrieval of archived LISTSERV messages

Dartmouth Merged SIGL

Special Interest Group Lists

Electronic Conferences

Directory of Scholarly Electronic Conferences, by Diane K. Kovacs

Brent Chapman creator of Majordomo MUD FAQ

frequently asked questions and answers about MUDs

MUD Lists

The Totally Unofficial List of Internet MUDs

MUD WWW Page Experimental MUD Page, information on MUDs MUDS

Virtual Spaces

MUDWHO via WWW

MUDWHO Access via WWW

MUD/Lysators Info about Lysators MUD’s, and MUD’s in general Free For All

an example of IW–allows user to place an entry in a list of WWW entries

BBS Home Page An experiment in interactive webbing IRC definition Internet Relay Chat Protocol, by Oikarinen and Reed IRC Info/Larsson IRC related documents and other sources of information, by Jonas Larsson IRC FAQ

frequently asked questions and answers list from newsgroup

IRC guide

Basic Guide to IRC

IRC documentation real-time, many-many text discussion divided into channels IRC via telnet using telnet to Undernet IRC via Unix Unix command for hosts with an installed IRC client ITR introduction Internet Talk Radio, by Carl Malamud ITR FAQ

Frequently Asked Questions on ITR (Usenet FAQ)

Mbone FAQ (text) Virtual Internet Backbone for Multicast IP Mbone FAQ Web Frequently Asked Questions (FAQ) on the Multicast Backbone (MBONE) Mbone Tools

lists suite of tools used with Mbone

USENET description What is Usenet?

USENET FAQs collection of frequently answer questions and answers about Usenet USENET documentation

special issue of Amateur Computerist Newsletter about Usenet

NNTP

Network News Transfer Protocol

Usenet Lists

news.groups FAQ directory

Search Usenet WAIS WAIS-based search for a Usenet and LISTSERV conferences Usenet filter

personalized netnews delivery service

Usenet filter

personalized netnews delivery service

Beginners group for people new to USENET Usenet reader nn, xrn)mand for hosts with newsreader, rn (other readers Usenet guide

An Introduction to Usenet News and the trn Newsreader, by Jon Bell

Usenet post via email post to newsgroup via email Usenet via WWW access newsgroups via WWW Usenet periodics list of Usenet periodic postings Usenet posters find when and from what host person with name FIRST LAST posted to Usenet IP

Internet Protocol

Kerberos

network authentication system for physically insecure networks

NFS Network File System will mount remote file systems across homogenous and heterogenous systems

PPP

Internet Standard for transmission of IP packets over serial lines

PPP FAQ

comp.protocols.ppp frequently wanted information

SLIP

Serial Line Internet Protocol

SNMP

Simple Network Management Protocol

SOCKS

A proxy server for IP hosts behind firewalls

TCP

Transmission Control Protocol

Winsock

accessing the Internet using tcp/ip applications under Microsoft Windows

ACC anonymous credit card (ACC) protocol, used for anonymous funds xfer and delivery on the Internet

Telecom stds

International Standards for telecommunication

ISO Protocols

newsgroup

OSI

Open Systems Interconnection protocols

URL guide

A Beginner’s Guide to URLs

URL syntax

Uniform Resource Locator Syntax

Internet Tools EARN guide to network resource tools Internet Tools NIR A status report on networked information retrieval tools and groups Net Mgt Tools Tools for Monitoring and Debugging TCP/IP Internets and Interconnected Devices PC Communications PC TCP/IP + NFS FAQ list, by Rawn Shah